# Wake Vision: A Large-scale, High-Quality Dataset and Benchmark Suite for TinyML Applications

### v2

@@ Appendix A Code Repository
The code used to generate the Wake Vision dataset is available at the following GitHub repository: https://github.com/harvard-edge/Wake_Vision
@@
Appendix B Flowchart of Bounding Box Filtering Process
@@
Appendix C Model Design Case Study
Basic test set performance can often misrepresent a model’s performance, given the typical domain shift between images scraped from the internet and real-world use cases. This issue is exacerbated when ML practitioners must trade off accuracy for model performance and size, which is necessary for TinyML use cases. A design decision might have seemingly little impact on the test accuracy but may destroy real-world performance depending on the deployment environment. For example, a person detection system may only operate in dark lighting conditions, but the test dataset has an insignificant number of dark samples; therefore the test accuracy will not reflect the real world accuracy.
@@ Appendix C Model Design Case Study
C.1 Image Size vs. Model Width Scaling
We train two series of MobileNetV2 models: one series that sweeps the input image size [64-256] and one that sweeps the width multiplier of a model [0.1-1.5]. We then benchmark these models on the Wake Vision test set as well as the far distance benchmark. We plot these results against the number of multiply accumulate (MAC) operations in the model as a proxy for on-device latency [2].
@@ C.2 Quantization
C.3 Grayscale
Converting a model’s input image channels to grayscale from RGB is a commonly employed optimization in the TinyML field [2] as it can substantially reduce a model’s memory consumption. We observed, however, that the grayscale optimization disproportionately impacts images on the brighter end of the spectrum as illustrated in Figure 6. This further demonstrates the importance of fine grained analysis, as some real world deployment environments might be far brighter than the average Wake Vision test sample.
@@ C.3 Grayscale
Appendix D Inducing Label Errors
@@
Appendix E Fine Grained Benchmark Images
@@
Appendix F Open Image Label Distribution
@@
Appendix G Manual Labeling
The validation and test sets were manually labeled through the crowd-sourced labeling platform Scale Rapid(https://scale.com/. Figure 8 shows a screenshot of the labeling menu. The labelers were instructed to label an image as a "person" if a person was present anywhere in the image and "no person" if no visible person was present. The labelers also indicated if the image contained a depiction of a person, which meant that sample was labeled "no person" in Wake Vision, or a "picture of a picture/reflection", which we used for metadata.
@@ Appendix H Dataset Access and Organization
Wake Vision is available online at the Harvard Dataverse, via TFDS, or Hugging Face Datasets. More information can be found on the https://wakevision.ai webpage.
The dataset is organized as a set of compressed tar files containing images, and a series of label CSVs. The label CSVs are organized such that the file name of the image is the identifying index. CSVs for all dataset splits include the person and depiction label. The Validation and Test label CSVs also have flags that denote a sample’s inclusion into a fine-grain benchmark set (e.g. Distance-Near). This structure makes it easy to access just the required data without requiring a full download. It also ensures the dataset can be easily updated as new versions are introduced.
@@ Appendix I Label Generation Details
Person Labels. The most straightforward Open Images label classes to label as person in Wake Vision are the "person" label and its subcategories (listed in appendix L). All of these are relabelled as persons in Wake Vision. These labels are present as both image-level labels and bounding box labels.
We furthermore inspect the image level label classes for synonyms and umbrella terms for all the person related labels. This search resulted in an additional six person related label classes. These person related label classes will only be used in the image level label configuration.
@@
Bounding Box Size. The VWW dataset only considered a Common Objects in Context (COCO) person to be a person if the bounding box around the person took up at least 5% of the image [5]. If the person took up less than 5% of the image, the image was excluded from the dataset. To make our dataset work as a plug in replacement for VWW, we adopt the same defaults in Wake Vision. For different requirements, users can change a configuration parameter in our open-source dataset creation code.
@@
To allow for both data-centric and model-centric improvements, we provide a standard model for data-centric contributions, and a standard pre-processing pipeline for model-centric contributions. Both types of contributions are expected to use accuracy as the primary metric for the overall test and validation set, and F1-score as the primary metric for the fine grained benchmarks introduced in section 4.
@@
The full Open Images v7 dataset is not hosted by the dataset authors. Rather it is provided as a collection of flickr image Uniform Resource Locators (URLs) and their associated labels. As an unfortunate result of this, the dataset is not static over time as image owners can delete their images from the flickr platform. As a result we were only able to download a subset of the original Open Images v7 dataset as shown in table 6.
@@
We consider the following Open Images v7 labels to be a person for the Wake Vision dataset:
@@
Images containing the following Open Images v7 labels and no other person related labels are excluded from the Wake Vision dataset:
@@
Appendix N Automatic Label Correction
@@
To further correct label errors in the final Wake Vision Validation and Test sets we employed the Scale AI platform to crowd-source manual label corrections. These label corrections are described in section 3.2 and deprecate the automatic label corrections described in this section.

### v3

@@ Appendix B Code Repository
The code used to generate the Wake Vision dataset is available at the following GitHub repository: https://anonymous.4open.science/r/Wake_Vision-DD19/README.md
@@
Appendix C Flowchart of Bounding Box Filtering Process
@@
Appendix D Model Design Case Study
Basic test set performance can often misrepresent a model’s performance, given the typical domain shift between images scraped from the internet and real-world use cases. This issue is exacerbated when ML practitioners must trade off accuracy for model performance and size, which is necessary for TinyML use cases. A design decision might have seemingly little impact on the test accuracy but may destroy real-world performance depending on the deployment environment. For example, a person detection system may only operate in dark lighting conditions, but the test dataset has an insignificant number of dark samples; therefore, the test accuracy will not reflect the real-world accuracy.
@@ Appendix D Model Design Case Study
D.1 Image Size vs. Model Width Scaling
We train two series of MobileNetV2 models: one series that sweeps the input image size [64-256] and one that sweeps the width multiplier of a model [0.1-1.5]. We then benchmark these models on the Wake Vision test set as well as the far distance benchmark. We plot these results against the number of multiply-accumulate (MAC) operations in the model as a proxy for on-device latency [2].
@@ D.2 Quantization
D.3 Grayscale
Converting a model’s input image channels to grayscale from RGB is a commonly employed optimization in the TinyML field [2] as it can substantially reduce a model’s memory consumption. We observed, however, that the grayscale optimization disproportionately impacts images on the brighter end of the spectrum as illustrated in Figure 8. This further demonstrates the importance of fine-grained analysis, as some real-world deployment environments might be far brighter than the average Wake Vision test sample.
@@ D.3 Grayscale
Appendix E Inducing Label Errors
@@
Appendix F Fine Grained Benchmark Images
@@
Appendix G Open Image Label Distribution
@@ Appendix H Dataset Access and Organization
Wake Vision is available online via TFDS and Hugging Face Datasets. We exclude direct links to the dataset during the review period to avoid breaking anonymity.
The dataset is organized as a set of compressed tar files containing images and a series of label CSVs. The label CSVs are organized such that the file name of the image is the identifying index. CSVs for all dataset splits include the person and depiction label. The Validation and Test label CSVs also have flags that denote a sample’s inclusion into a fine-grain benchmark set (e.g. Distance-Near). This structure makes it easy to access just the required data without requiring a full download. It also ensures the dataset can be easily updated as new versions are introduced.
@@ Appendix I Label Generation Details
Person Labels. The most straightforward Open Images label classes to label as a person in Wake Vision are the "person" label and its subcategories (Configurable, but defaults listed in Appendix L). All of these are relabelled as persons in Wake Vision. These labels are present as both image-level labels and bounding box labels.
We furthermore inspect the image-level label classes for synonyms and umbrella terms for all the person related labels. This search resulted in an additional six person related label classes. These person related label classes will only be used in the image level label configuration.
Lastly, the image level person labels in Open Images have an associated confidence score. This score ranges from 0-10. Labels that have been verified by humans to be absent from an image have a score of 0, while labels that have been verified present by humans have a score of 10. Purely machine-generated labels have a fractional confidence score that is generally $>=5$ [21, 19]. By default, we only respect labels that have a minimum confidence of 7. Labels below this threshold are ignored.
@@
Bounding Box Size. The VWW dataset only considered a Common Objects in Context (COCO) person to be a person if the bounding box around the person took up at least 0.5% of the image [5]. If the person took up less than 0.5% of the image, the image was excluded from the dataset. For Wake Vision, we observed a decrease in performance at the 0.5% limit. We theorize that the cause of this decrease is that a person becomes indistinguishable at such a low percentage of already small images. Therefore we adopted a 5% threshold for the Wake Vision dataset instead. For different requirements, users can change a configuration parameter in our open-source dataset creation code.
@@
To allow for both data-centric and model-centric improvements, we provide a standard model for data-centric contributions, and a standard pre-processing pipeline for model-centric contributions. Both types of contributions are expected to use accuracy as the primary metric for the overall test and validation set, and F1-score as the primary metric for the fine grained benchmarks introduced in Sec. 3.3.
@@
The full Open Images v7 dataset is not hosted by the dataset authors. Rather it is provided as a collection of flickr image Uniform Resource Locators (URLs) and their associated labels. As an unfortunate result of this, the dataset is not static over time as image owners can delete their images from the flickr platform. As a result we were only able to download a subset of the original Open Images v7 dataset as shown in Tab. 7.
@@
In our default configuration, we consider the following Open Images v7 labels to be a person for the Wake Vision dataset:
@@
The following labels are considered body parts and are labelled according to the body part configuration.:
@@
Appendix N Manual Label Correction
We use the Scale Rapid tool from Scale AI to relabel the Wake Vision validation and test sets. Figure 10 shows a screenshot of the labeling menu presented to the human labelers. The labelers were instructed to label an image as a "person" if a person was present anywhere in the image and "no person" if no visible person was present. The labelers also indicated if the image contained a depiction of a person or a "picture of a picture/reflection", which we use for metadata.
Appendix O Automatic Label Correction
@@
To further correct label errors in the final Wake Vision Validation and Test sets we employed the Scale AI platform to crowd-source manual label corrections. These label corrections are described in Sec. 3.2 and deprecate the automatic label corrections described in this section.
Appendix P Author Statement
The authors of this paper hereby confirm that we bear all responsibility in case of violation of rights, etc, and confirm the data license to be CC BY 4.0 for our labels, and that all images in the dataset have been uploaded to Flickr with a CC BY 2.0 license. Note: while the Open Images authors tried to identify images that are licensed under a Creative Commons Attribution license, we make no representations or warranties regarding the license status of each image and users should verify the license for each image themselves.
Appendix Q Hosting, License and Maintenance Plan
The dataset is hosted on HuggingFace Datasets. Additionally, the dataset is available through TensorFlow datasets and can be regenerated from Open Images using our open-source filtering code. The Edge AI Foundation will assume responsibility for hosting and maintaining the dataset long-term.
Wake Vision’s labels and metadata are licensed under a CC BY 4.0 license. Images in the dataset have all been uploaded to Flickr with a CC BY 2.0 license. Note: while the Open Images authors tried to identify images that are licensed under a Creative Commons Attribution license, we make no representations or warranties regarding the license status of each image and users should verify the license for each image themselves.
Appendix R DOI and Croissant URLs
The DOI and URLs will be provided after the review period to avoid breaking anonymity.
Appendix S Datasheet for Wake Vision
This document is based on Datasheets for Datasets by Gebru et al. [12]. Please see the most updated version here.
{mdframed}
[linecolor=black]
MOTIVATION
For what purpose was the dataset created? Was there a specific task in mind? Was there a specific gap that needed to be filled? Please provide a description.
Wake Vision was created as a prototype for a new generation of datasets for TinyML. Where previous TinyML datasets were small due to limited resources, Wake Vision is a production grade dataset consisting of almost 6 million images and with an extensively cleaned validation and test set. Wake Vision specifically targets the binary person classification task. Such a large dataset enables TinyML researchers to explore questions previously impossible. For example, the size of the Wake Vision dataset enables new research into the tradeoff that often exist between dataset size and sample quality.
Who created this dataset (e.g., which team, research group) and on behalf of which entity (e.g., company, institution, organization)?
These details will be provided after the review period to avoid breaking anonymity.
What support was needed to make this dataset? (e.g.who funded the creation of the dataset? If there is an associated grant, provide the name of the grantor and the grant name and number, or if it was supported by a company or government agency, give those details.)
These details will be provided after the review period to avoid breaking anonymity.
Any other comments?
No additional comments
{mdframed}
[linecolor=black]
COMPOSITION
What do the instances that comprise the dataset represent (e.g., documents, photos, people, countries)? Are there multiple types of instances (e.g., movies, users, and ratings; people and interactions between them; nodes and edges)? Please provide a description.
The dataset is composed of images, labels, and metadata. The images are sourced from Open Images. The labels are binary, with ’1’ indicating there is a person somewhere in the image and ’0’ indicating that there is not a person in the image. The metadata is used to form the benchmark suite where we can evaluate models in specific challenging settings. Each image has a binary variable for each fine-grained benchmark where a ’1’ indicates that the image belongs to the fine-grained benchmark set.
How many instances are there in total (of each type, if appropriate)?
There are over 6 million labeled images in the dataset.
Does the dataset contain all possible instances or is it a sample (not necessarily random) of instances from a larger set? If the dataset is a sample, then what is the larger set? Is the sample representative of the larger set (e.g., geographic coverage)? If so, please describe how this representativeness was validated/verified. If it is not representative of the larger set, please describe why not (e.g., to cover a more diverse range of instances, because instances were withheld or unavailable).
The dataset uses a large subset of the Open Images dataset (6 million out of a possible 9 million images). We use a subset in order to have a balanced number of person samples and non-person samples. We also filter the dataset for quality. The filtering removes subjects that are too far and outside of a center crop. Open Images sources images from Flickr that are licensed as CC-BY 2.0. This is likely not a fully representative sample of the overall world population and likely skews heavily towards a European or North American population. See the Open Images paper [21] for more information.
What data does each instance consist of? “Raw” data (e.g., unprocessed text or images) or features? In either case, please provide a description.
The dataset is composed of images, labels, and metadata.
Is there a label or target associated with each instance? If so, please provide a description.
Yes, every image has a binary label to indicate if a person is visible in the image. Images in the some dataset splits include more metadata labels, e.g., about distance of subject or whether the subject is a depiction.
Is any information missing from individual instances? If so, please provide a description, explaining why this information is missing (e.g., because it was unavailable). This does not include intentionally removed information, but might include, e.g., redacted text.
The training set does not contain the metadata used for the fine grained benchmark suite. The dataset contains two training sets, one of higher quality and one larger. The higher quality dataset and other sets contain information about depictions and body parts. The large training set does not contain this information.
Are relationships between individual instances made explicit (e.g., users’ movie ratings, social network links)? If so, please describe how these relationships are made explicit.
No
Are there recommended data splits (e.g., training, development/validation, testing)? If so, please provide a description of these splits, explaining the rationale behind them.
Yes there are separate splits for training, validation, and testing. These splits are generated from equivalent splits in Open Images. The dataset contains two training splits, one focusing on quality of samples and one on number of samples.
Are there any errors, sources of noise, or redundancies in the dataset? If so, please provide a description.
Yes there are label errors in the dataset due to the labeling process. We discuss this at length in the paper.
Is the dataset self-contained, or does it link to or otherwise rely on external resources (e.g., websites, tweets, other datasets)? If it links to or relies on external resources, a) are there guarantees that they will exist, and remain constant, over time; b) are there official archival versions of the complete dataset (i.e., including the external resources as they existed at the time the dataset was created); c) are there any restrictions (e.g., licenses, fees) associated with any of the external resources that might apply to a future user? Please provide descriptions of all external resources and any restrictions associated with them, as well as links or other access points, as appropriate.
The dataset is self-contained. We re-host the images, labels, and metadata. The images may be subject to different licenses than our labels and metadata. All images have originally been hosted on Flickr under a permissive CC BY license.
Does the dataset contain data that might be considered confidential (e.g., data that is protected by legal privilege or by doctor-patient confidentiality, data that includes the content of individuals’ non-public communications)? If so, please provide a description.
No, the images were posted online with the CC-BY 2.0 license. It is possible an image was posted without the individual’s consent. While we tried to identify images that are licensed under a Creative Commons Attribution license, we make no representations or warranties regarding the license status of each image, and a user should verify the license for each image themselves.
Does the dataset contain data that, if viewed directly, might be offensive, insulting, threatening, or might otherwise cause anxiety? If so, please describe why.
No, not to our knowledge. Due to the size of the dataset we have not had the resources to manually verify this.
Does the dataset relate to people? If not, you may skip the remaining questions in this section.
Yes, the dataset contains images of people.
Does the dataset identify any subpopulations (e.g., by age, gender)? If so, please describe how these subpopulations are identified and provide a description of their respective distributions within the dataset.
Yes the benchmark suite contains subsets that focus on perceived age and gender. These labels have been sourced from the More Inclusive Annotations of People [34]. The purpose of these benchmarks are to test the fairness of a trained model across these subpopulation and we do not condone using these labels to train a model. The size of these sets is listed in Table 5.
Is it possible to identify individuals (i.e., one or more natural persons), either directly or indirectly (i.e., in combination with other data) from the dataset? If so, please describe how.
It could be possible to identify a person by their face or other visible characteristics in the image. We do not share any other personal information besides the images.
Does the dataset contain data that might be considered sensitive in any way (e.g., data that reveals racial or ethnic origins, sexual orientations, religious beliefs, political opinions or union memberships, or locations; financial or health data; biometric or genetic data; forms of government identification, such as social security numbers; criminal history)? If so, please provide a description.
No, not to our knowledge. Some personal or sensitive information may be visible in the images. However, the images were shared publicly under a CC-BY 2.0 license. We do not include any personal information beyond the images.
Any other comments?
The images used were published publicly under a CC-BY 2.0 license and are already used in the Open Images [19] dataset and its derivatives. The authors of Open images tried to identify images that are licensed under a Creative Commons Attribution license but make no representations or warranties regarding the license status of each image and a user should verify the license for each image themselves.
{mdframed}
[linecolor=black]
COLLECTION
How was the data associated with each instance acquired? Was the data directly observable (e.g., raw text, movie ratings), reported by subjects (e.g., survey responses), or indirectly inferred/derived from other data (e.g., part-of-speech tags, model-based guesses for age or language)? If data was reported by subjects or indirectly inferred/derived from other data, was the data validated/verified? If so, please describe how.
The dataset is derived from Open Images [19], which in turn acquired its images from Flickr. Section 3.1 describes the process.
Over what timeframe was the data collected? Does this timeframe match the creation timeframe of the data associated with the instances (e.g., recent crawl of old news articles)? If not, please describe the timeframe in which the data associated with the instances was created. Finally, list when the dataset was first published.
The original data collection was conducted in Noember of 2015. The Open Images paper [21] details the image acquisition process.
What mechanisms or procedures were used to collect the data (e.g., hardware apparatus or sensor, manual human curation, software program, software API)? How were these mechanisms or procedures validated?
The images were sources from Flickr in an automated pipeline. The original labeling pipeline was a combination of automated labeling and manual verification and labeling. The Open Images paper details this process [21]. We detail our process for relabeling and filtering the dataset in Section 3.1.
What was the resource cost of collecting the data? (e.g. what were the required computational resources, and the associated financial costs, and energy consumption - estimate the carbon footprint. See Strubell et al.[35] for approaches in this area.)
We used cloud TPU credits provided by the Google Cloud TRC program. The primary source of required resources was model training to evaluate the dataset and the storage and bandwidth required to process and upload the dataset to hosting locations. The total size of the dataset is approximately 2 TB. To the authors knowledge, there is no public information about the power draw of TPUs. This matches the conclusions drawn by Strubell et al.[35]. This makes it impossible to disclose the energy consumption and estimated carbon footprint of the project.
We used paid crowdsource labelers to relabel the validation and test sets. The total cost of labeling was approximately \$7,000.
If the dataset is a sample from a larger set, what was the sampling strategy (e.g., deterministic, probabilistic with specific sampling probabilities)?
The dataset is a large subset of the total Open Images dataset. The subset is generated by a deterministic filtering process described in Section 3.1.
Who was involved in the data collection process (e.g., students, crowdworkers, contractors) and how were they compensated (e.g., how much were crowdworkers paid)?
We used paid crowdsource labelers to relabel the validation and test sets. The total cost of labeling was approximately \$7,000. We used the Scale Rapid labeling service. A total of 70,898 images were labeled at a cost of \$0.1 per image and \$0.033 per individual label. Appendix 3.2 gives additional detail.
Were any ethical review processes conducted (e.g., by an institutional review board)? If so, please provide a description of these review processes, including the outcomes, as well as a link or other access point to any supporting documentation.
No
Does the dataset relate to people? If not, you may skip the remainder of the questions in this section.
Yes the dataset contains images of people.
Did you collect the data from the individuals in question directly, or obtain it via third parties or other sources (e.g., websites)?
We obtained the images via the Open Images dataset [21] which originally sourced its images from Flickr [10].
Were the individuals in question notified about the data collection? If so, please describe (or show with screenshots or other information) how notice was provided, and provide a link or other access point to, or otherwise reproduce, the exact language of the notification itself.
No the individuals were not notified to our knowledge. The images were posted publicly under the CC-BY 2.0 license.
Did the individuals in question consent to the collection and use of their data? If so, please describe (or show with screenshots or other information) how consent was requested and provided, and provide a link or other access point to, or otherwise reproduce, the exact language to which the individuals consented.
Yes, the images were posted publicly under the CC-BY 2.0 license. It is possible that the images were uploaded by a third party without the consent of the individual in the image.
If consent was obtained, were the consenting individuals provided with a mechanism to revoke their consent in the future or for certain uses? If so, please provide a description, as well as a link or other access point to the mechanism (if appropriate)
No we do not have a mechanism for an individual the revoke their consent. As with Open Images, we make no representations or warranties regarding the license status of each image, and a user should verify the license for each image themselves via the Flickr link in the metadata. Note that it is impossible to revoke a CC BY license after publication [23].
Has an analysis of the potential impact of the dataset and its use on data subjects (e.g., a data protection impact analysis)been conducted? If so, please provide a description of this analysis, including the outcomes, as well as a link or other access point to any supporting documentation.
No formal analysis has been conducted. We describe the ethical considerations of the dataset in Section 5.
Any other comments?
No additional comments.
{mdframed}
[linecolor=black]
PREPROCESSING / CLEANING / LABELING
Was any preprocessing/cleaning/labeling of the data done(e.g.,discretization or bucketing, tokenization, part-of-speech tagging, SIFT feature extraction, removal of instances, processing of missing values)? If so, please provide a description. If not, you may skip the remainder of the questions in this section.
Yes we filtered the data as shown in Figure 6.
Was the “raw” data saved in addition to the preprocessed/cleaned/labeled data (e.g., to support unanticipated future uses)? If so, please provide a link or other access point to the “raw” data.
The "raw" data is accessible in the Open Images dataset.
Is the software used to preprocess/clean/label the instances available? If so, please provide a link or other access point.
Yes, we open-source our filtering code. https://anonymous.4open.science/r/Wake_Vision-DD19/README.md
Any other comments?
No additional comments
{mdframed}
[linecolor=black]
USES
Has the dataset been used for any tasks already? If so, please provide a description.
The dataset has been used for the experiments reported in the paper.
Is there a repository that links to any or all papers or systems that use the dataset? If so, please provide a link or other access point.
The usage of the dataset can be tracked via the citations of this paper. We plan to host competitions using the dataset and will track participants.
What (other) tasks could the dataset be used for?
The dataset can be used as a benchmark for TinyML models and as the basis for data-centric AI research given the prevalence of label errors in the training set and the cleaned nature of the validation and test sets.
Is there anything about the composition of the dataset or the way it was collected and preprocessed/cleaned/labeled that might impact future uses? For example, is there anything that a future user might need to know to avoid uses that could result in unfair treatment of individuals or groups (e.g., stereotyping, quality of service issues) or other undesirable harms (e.g., financial harms, legal risks) If so, please provide a description. Is there anything a future user could do to mitigate these undesirable harms?
Yes, the estimated error rates are reported in Table 2. We also discuss the high likelihood of false negatives in Wake Vision’s training sets in Section 3.2. These errors are due to the labeling and filtering process. We mitigate the impact of these errors by manually labeling the validation and test sets as well as study their impact for TinyML models in Section 4.4.
Are there tasks for which the dataset should not be used? If so, please provide a description.
The perceived age and gender benchmarks should not be used to train a gender or age classifier.
Any other comments?
No additional comments
{mdframed}
[linecolor=black]
DISTRIBUTION
Will the dataset be distributed to third parties outside of the entity (e.g., company, institution, organization) on behalf of which the dataset was created? If so, please provide a description.
The dataset is hosted on Harvard Dataverse and HuggingFace Datasets. Additionally, the dataset is available through TensorFlow datasets and can be regenerated from Open Images with our open source filtering code. We plan for the TinyML foundation (https://www.tinyml.org/) to assume responsibility for hosting and maintaining the dataset long-term.
How will the dataset will be distributed (e.g., tarball on website, API, GitHub)? Does the dataset have a digital object identifier (DOI)?
Yes, the dataset is hosted on Harvard Dataverse and HuggingFace Datasets. Additionally, the dataset is available through TensorFlow datasets and can be regenerated from Open Images with our open source filtering code.
When will the dataset be distributed?
The Dataset is publicly accessible at the time of writing.
Will the dataset be distributed under a copyright or other intellectual property (IP) license, and/or under applicable terms of use (ToU)? If so, please describe this license and/or ToU, and provide a link or other access point to, or otherwise reproduce, any relevant licensing terms or ToU, as well as any fees associated with these restrictions.
The dataset is published under the CC-BY 4.0 license.
Have any third parties imposed IP-based or other restrictions on the data associated with the instances? If so, please describe these restrictions, and provide a link or other access point to, or otherwise reproduce, any relevant licensing terms, as well as any fees associated with these restrictions.
No
Do any export controls or other regulatory restrictions apply to the dataset or to individual instances? If so, please describe these restrictions, and provide a link or other access point to, or otherwise reproduce, any supporting documentation.
No
Any other comments?
No additional comments.
{mdframed}
[linecolor=black]
MAINTENANCE
Who is supporting/hosting/maintaining the dataset?
The dataset is hosted on HuggingFace Datasets. Additionally, the dataset is available through TensorFlow datasets and can be regenerated from Open Images with our open-source filtering code. The Edge AI Foundation will assume responsibility for hosting and maintaining the dataset long-term.
How can the owner/curator/manager of the dataset be contacted (e.g., email address)?
Removed during review to maintain anonymity.
Is there an erratum? If so, please provide a link or other access point.
No
Will the dataset be updated (e.g., to correct labeling errors, add new instances, delete instances)? If so, please describe how often, by whom, and how updates will be communicated to users (e.g., mailing list, GitHub)?
The dataset may be updated if any problematic instances are discovered or the label errors are corrected. In this case, we will issue a new version of the dataset to ensure past results are contextualized under the version of the dataset that was used.
If the dataset relates to people, are there applicable limits on the retention of the data associated with the instances (e.g., were individuals in question told that their data would be retained for a fixed period of time and then deleted)? If so, please describe these limits and explain how they will be enforced.
There is no limit.
Will older versions of the dataset continue to be supported/hosted/maintained? If so, please describe how. If not, please describe how its obsolescence will be communicated to users.
Yes, we will continue to host older versions of the dataset. Details will be provided after review to preserve anonymity.
If others want to extend/augment/build on/contribute to the dataset, is there a mechanism for them to do so? If so, please provide a description. Will these contributions be validated/verified? If so, please describe how. If not, why not? Is there a process for communicating/distributing these contributions to other users? If so, please provide a description.
Yes, the dataset is licensed permissively, and the code used to create Wake Vision is open source.
Any other comments?
No additional comments